Hydrogen sulfide gas is combusted with oxygen gas to produce sulfur dioxide gas and water vapour. If there is 48.4 L of oxygen available, what volume of hydrogen sulfide gas may be combusted if the pressure is held constant at 105 kPa and the temperature is held constant at 190°C?


1
Expert's answer
2021-05-18T07:52:48-0400

2H2S + 3O2 → 2SO2 + 2H2O

V(O2) = 48.4 L

p = 105 kPa = 1.036 atm

T = 190 + 273 = 463 K

Ideal gas law:

pV = nRT

"n = \frac{pV}{RT}"

R = 0.08206 L×atm/mol×K

n(O2) "= \frac{1.036 \times 48.4}{0.08206 \times 463}=1.319 \; mol"

According to the reaction:

n(H2S) = "\frac{2}{3}" n(O2) = "\frac{2}{3} \times 1.319 = 0.8798 \;mol"

"V = \frac{nRT}{p} \\

V(H_2S) = \frac{0.8798 \times 0.08206 \times 463}{1.036}=32.26 \;L"

Answer: 32.26 L